### v3

@@ 6 On the optimality of the regularity
where $\kappa_{\epsilon}>0$ is a positive constant independent of $j$. This constant is simply the probability that a standard Brownian motion stays inside the interval $\epsilon,\epsilon]$ throughout the time interval $[0,1]$. This is a classical fact about the distribution of the maximum of Brownian motion, see e.g. [Res92, Section 6.7].
For $P_{2}$, we have
@@
$P_{2}$ $=\mathbb{P}\left(c-\epsilon\sqrt{|I_{j}|}\leq W_{T-\delta 2^{-j}}\leq c+\epsilon\sqrt{|I_{j}|}\right)$
$=\frac{1}{\sqrt{2\pi(T-\delta 2^{-j})}}\int_{c-\epsilon\sqrt{|I_{j}|}}^{c+\epsilon\sqrt{|I_{j}|}}\exp\left({-\frac{1}{2}\frac{u^{2}}{T-\delta 2^{-j}}}\right)\textnormal{d}{u}$
Since $c>0$ and $\epsilon>0$ are given, there is $j_{0}\in\mathbb{N}$ large enough such that $c-\epsilon\sqrt{|I_{j}|}>0$ for all $j\geq j_{0}$. Therefore,
@@
$P_{2}\geq\frac{2\epsilon\sqrt{|I_{j}|}}{\sqrt{2\pi T}}\exp\left(-\frac{1}{2}\frac{\left(c+\epsilon T\right)^{2}}{T-\delta}\right)=\tilde{\kappa}_{c,\epsilon,T,\delta}\sqrt{|I_{j}|}$ (125)
for some positive constant $\tilde{\kappa}_{c,\epsilon,T,\delta}$ independent of $j$. Combining (124) and (125), we obtain the desired result. ∎

### v5

@@ 6 On the optimality of the regularity
and $\widetilde{W}$ is a standard Brownian motion. Here $\kappa_{\varepsilon}>0$ is a positive constant independent of $N$. This constant is simply the probability that a standard Brownian motion remains inside $\varepsilon,\varepsilon]$ throughout the time interval $[0,1]$. This is a classical fact concerning the distribution of the maximum of Brownian motion, see, for instance, [RES92, Section 6.7]. In particular, $q_{N}>0$.
Assume now that $q_{i+1}>0$ and let $x\in\tilde{I}_{i}$. We show that $q_{i}>0$. In more detail,
$\psi_{i}(x)$ $=\mathbb{E}\!\left(\mathbf{1}_{B_{i}}\,\psi_{i+1}(W_{t_{i+1}})\,\middle|\,W_{t_{i}}=x\right)$
$\geq q_{i+1}\,\mathbb{P}\!\left(B_{i}\cap\Big\{|W_{t_{i+1}}-c|\leq\tfrac{\varepsilon}{4\sqrt{2}}\sqrt{|I_{i}|}\Big\}\,\middle|\,W_{t_{i}}=x\right).$ (129)
Here we used the definition of $q_{i+1}$ together with the fact that $|I_{i+1}|=\tfrac{1}{2}|I_{i}|$. Indeed,
$|W_{t_{i+1}}-c|\leq\tfrac{\varepsilon}{4\sqrt{2}}\sqrt{|I_{i}|}=\tfrac{\varepsilon}{4}\sqrt{|I_{i+1}|}$
implies that $W_{t_{i+1}}\in\tilde{I}_{i+1}$, and therefore
$\psi_{i+1}(W_{t_{i+1}})\geq q_{i+1}.$
To bound the probability appearing in (129), we apply a rescaling argument analogous to that used in (128). Conditioning on $W_{t_{i}}=x$ and using the scaling property of Brownian motion, we obtain
$\mathbb{P}\!\left(B_{i}\cap\Big\{|W_{t_{i+1}}-c|\leq\tfrac{\varepsilon}{4\sqrt{2}}\sqrt{|I_{i}|}\Big\}\,\middle|\,W_{t_{i}}=x\right)$
$=\mathbb{P}\!\left(\left\{\sup_{s\in[0,1]}|\widetilde{W}_{s}|\leq\varepsilon\right\}\cap\left\{|\widetilde{W}_{1}-m|\leq\tfrac{\varepsilon}{4\sqrt{2}}\right\}\right),$ (130)
where $m=\frac{c-x}{\sqrt{|I_{i}|}}$. Since $x\in\tilde{I}_{i}$, we have
$|m|=\frac{|c-x|}{\sqrt{|I_{i}|}}\leq\frac{\varepsilon}{4}.$
Consequently,
$\Big[-\tfrac{\varepsilon}{4\sqrt{2}}+m,\;\tfrac{\varepsilon}{4\sqrt{2}}+m\Big]\subset(-\varepsilon,\varepsilon).$
Therefore,
$\mathbb{P}\!\left(\left\{\sup_{s\in[0,1]}|\widetilde{W}_{s}|\leq\varepsilon\right\}\cap\left\{|\widetilde{W}_{1}-m|\leq\tfrac{\varepsilon}{4\sqrt{2}}\right\}\right)$
@@
$=\mathbb{P}\big(\sup_{s\in[0,1]}|\widetilde{W}_{s}|\leq\varepsilon\big)\,\mathbb{P}\!\left(\widetilde{W}_{1}\in\Big[-\tfrac{\varepsilon}{4\sqrt{2}}+m,\,\tfrac{\varepsilon}{4\sqrt{2}}+m\Big]\,\middle|\,\sup_{s\in[0,1]}|\widetilde{W}_{s}|\leq\varepsilon\right).$ (131)
Since the conditional law of $\widetilde{W}_{1}$ given $\{\sup_{s\in[0,1]}|\widetilde{W}_{s}|\leq\varepsilon\}$ admits a continuous and strictly positive density on $(-\varepsilon,\varepsilon)$, and since all admissible intervals above are contained in a fixed compact subset of $(-\varepsilon,\varepsilon)$, there exists a constant $c_{\varepsilon}>0$, depending only on $\varepsilon$, such that
$\mathbb{P}\!\left(\widetilde{W}_{1}\in\Big[-\tfrac{\varepsilon}{4\sqrt{2}}+m,\,\tfrac{\varepsilon}{4\sqrt{2}}+m\Big]\,\middle|\,\sup_{s\in[0,1]}|\widetilde{W}_{s}|\leq\varepsilon\right)\geq c_{\varepsilon},$
uniformly in $m$. Hence,
$\mathbb{P}\left(\left\{\sup_{s\in[0,1]}|\widetilde{W}_{s}|\leq\varepsilon\right\}\cap\left\{|\widetilde{W}_{1}-m|\leq\tfrac{\varepsilon}{4\sqrt{2}}\right\}\right)\geq\kappa_{\varepsilon}\,c_{\varepsilon}>0.$
By the induction hypothesis $q_{i+1}>0$, so it follows that
$\psi_{i}(x)\geq q_{i+1}\kappa_{\varepsilon}\,c_{\varepsilon}>0\qquad\text{for all }x\in\tilde{I}_{i}.$
Since $q_{i}:=\inf_{x\in\tilde{I}_{i}}\psi_{i}(x)$, this proves that $q_{i}>0$. Therefore, $q_{i}>0$ for all $i\in\{j_{0},\ldots,N\}$, and
$q_{i}\geq q_{N}(\kappa_{\varepsilon}c_{\varepsilon})^{N-i}.$
Finally, by the law of total probability and the definition of the regular conditional probability $\psi_{j_{0}}$, we obtain
$\mathbb{P}\!\left(\bigcap_{k=j_{0}}^{N}A_{k}\right)$ $=\mathbb{E}\!\left(\psi_{j_{0}}(W_{t_{j_{0}}})\right)$
$\geq\mathbb{E}\!\left(\psi_{j_{0}}(W_{t_{j_{0}}})\mathbf{1}_{\{W_{t_{j_{0}}}\in\tilde{I}_{j_{0}}\}}\right)$
@@
$\geq q_{N}(\kappa_{\varepsilon}c_{\varepsilon})^{N-j_{0}}\mathbb{P}\!\left(|W_{t_{j_{0}}}-c|\leq\tfrac{\varepsilon\sqrt{\delta\,2^{-j_{0}-1}}}{4}\right)$
$\geq q_{N}(\kappa_{\varepsilon}c_{\varepsilon})^{N-j_{0}}\tilde{\kappa}_{c,\varepsilon,T,\delta}\,\sqrt{|I_{j_{0}}|}>0,$
where $\tilde{\kappa}_{c,\varepsilon,T,\delta}>0$ is a constant depending only on the parameters $c$, $\varepsilon$, $T$, and $\delta$. This concludes the proof. ∎
@@ Acknowledgements
The authors are grateful to the anonymous referees for their careful reading of the manuscript and for their valuable comments and suggestions, which helped improve the quality and clarity of the paper.